OpenStudy (anonymous):

Eric and his crew have 50 trees to plant. In 8 hours they planted 40 trees. If they keep working at the same rate tomorrow, ho long will it take to finish planting the trees?

OpenStudy (anonymous):

8 hrs-40 trees x hrs-10 trees

OpenStudy (anonymous):

i divided 8 and 40 i got 2.50 hours

OpenStudy (anonymous):

x=8*10/40=80/40=2 hrs

OpenStudy (anonymous):

so it will take them 2 hours to plant 10 more trees.

OpenStudy (anonymous):

but what about the 50 mins?

OpenStudy (smarty):

.50 hours means half an hour.

OpenStudy (anonymous):

They have 50 trees to plant, they planted 40 in 8 hours, they have 10 trees to go.

OpenStudy (anonymous):

Make a simple proportion.